The length of a rectangle is twice its width. If the perimeter of the rectangle is 30ft, find its area
Answer by stanbon(57282)

The length of a rectangle is twice its width. If the perimeter of the rectangle is 30ft, find its area
---------------------------------
Perimeter = 2(L+W) = 30
L + W= 15
--------------------------
But L = 2W, so substitute to get:
2W + W = 15
3W = 15
W = 5
L = 2W = 10
-----------------------
Area = LW = 10*5 = 50 sq. ft.
================================
Cheers,
Stan H.